OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. 4x2 + ___x + 49

OpenStudy (anonymous):

for this, i have to put the exact answer in the answer box or i get it marked as wrong. can anyone give me the explaination for this along with the answer clearly marked as the answer ?

OpenStudy (anonymous):

well for a perfect square, the bx(what you are trying to find) multipies together the get c)in this case 49),

OpenStudy (anonymous):

uhm okk

OpenStudy (anonymous):

this is the form of ax^2+bx+c, correct?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

okay, so with knowing that we know that a=4, c=49, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so b is going to be the square root of C.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what would your answer be?

OpenStudy (anonymous):

idk.i have to put the exact answer in or its marked wrong so i dont trust my own judgement.

OpenStudy (anonymous):

well what do you think it is and ill tell you.

OpenStudy (anonymous):

what is the square root of 49?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

okay, now since there is a leading coefficient in frontof the x^2 what does that tell you?

OpenStudy (anonymous):

i dont have a clue. this is my weakest subject

OpenStudy (anonymous):

okay, well with that 4 there, it makes it a little more complicated. @hartnn can probably explain it better.

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

ok, if the leading co-efficient is not =1, you make it equal to 1 by dividing by 4, say you missing term is 'm' so, x^2 + (m/4)x +49/4 got this ?

OpenStudy (anonymous):

kinda

hartnn (hartnn):

so now (last term)^2 = 49/4 then last term =... ?

OpenStudy (anonymous):

ok u lost me

hartnn (hartnn):

if thats a perfect square, then its in the form, \((a+b)^2 =(a^2+2ab+b^2)\) right ? see the last term is b^2, to get 'b' you take square root of last term. here last term = 49/4 so what is square root of 49/4 ??

OpenStudy (anonymous):

7

hartnn (hartnn):

7 is the square root of 49 i asked square root of 49/4

OpenStudy (anonymous):

oh 3.5

hartnn (hartnn):

yeah.. keep it as 7/2 only. thats your 'b' to get the middle term m/4 , just multiply this by 2 , 7/2 *2 = ... ?